Please answer in the empty box
You are considering investing in Dakotas Security Services. You have been able to locate the following information on the firm: Total assets are $33.4 million, accounts receivable are $4.54 million, ACP is 25 days, net income is $4.80 million, and debt-to-equity is 12 times. All sales are on credit. Dakotas is considering loosening its credit policy such that ACP will increase to 30 days. The change is expected to increase credit sales by 5 percent. Any change in accounts receivable will be offset with a change in debt. No other balance sheet changes are expected. Dakotas profit margin will remain unchanged. How will this change in accounts receivable policy affect Dakotas net income, total asset turnover, equity multiplier, ROA, and ROE? Do not round intermediate calculations. Enter your answer in millions of dollars. Round your answers to 2 decimal places. Use 365 days a year.) Net income Total asset tumover Equity multiplier ROA ROE times times
Answer #1

ACP=(365*acc rec)/sales
oldsales=(365*4.54*10^6)/25=66284000
New sales=66284000*(1+5%)=69598200
New ACP=(365* NEw Acc)/New sales
New Acc rec=(30*69598200)/365=5720400
New Assets=(old assets+New acc rec-old acc rec)
=(33.4*10^6)+(5720400-(4.54*10^6))=34580400
Old profit margin=(net income/sales)*100
=(4.8*10^6)/66284000
=7.24%
New net income=7.24%*69598200=5.04mn
Total asset turnover= new sales/new assets
=(69598200/34580400)=2.01
Equity Multiplier=Assets/equity
Old debt to equity=1.2
old deb+old equity=33.4mn
Old debt=18.22mn
New debt is old debt+increase in acc rec
New debt=18218181.82+(5720400-(4.54*10^6))=19398581.82
New equity= new assets-new debt=34580400-19398581.82
=15181818.18
Equity multiplier=34580400/15181818.18=2.28
ROA=net income/assets
=(5.04*10^6)/34580400=14.57%
ROE=Net income/equity
=(5.04*10^6)/15181818.18=33.20%